OpenStudy (anonymous):

25 over 25 as a decimal step by step please

OpenStudy (anonymous):

1, no decimal needed...

OpenStudy (anonymous):

25 is a multiple of it self so, 25/25 reduces to 1/1 thus 25/25 = 1

OpenStudy (anonymous):

as a decimal the answer key says .96%

OpenStudy (anonymous):

Something divided by itself is always 1, is there more to the problem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(well 0/0 and \[\infty/\infty\] are a couple exceptions where that rule goes crazy)

OpenStudy (anonymous):

24over25 as a percent?

OpenStudy (anonymous):

how do they get 96%?

OpenStudy (anonymous):

There ya go, 24/25 would make sence for .96 Think of it as (24/25)*4/100

OpenStudy (anonymous):

sense*

OpenStudy (anonymous):

24 times 4? 25 times 4=?

OpenStudy (anonymous):

right you are trying to make it into a similar number to 96/100=.96%

OpenStudy (anonymous):

so to get 25 to equal 100 you need to take 25*4. Thus you would also need to take 24*4 to get the correct numerator.

OpenStudy (anonymous):

and a number divided by 100 is easily made into a percentage just by shifting the decimal place to places to the left